Shutdown asml SNF 2010-06-07 06:55:05: Scheduled Downtime 6/7

gsosa at snf.stanford.edu gsosa at snf.stanford.edu
Mon Jun 7 13:49:42 PDT 2010


Powered up tool after chilled water system was restored. Initialized hardware/software- OK. Re-started and warmed up UV lamp- Uniformity OK at 1.0%. Exposed hotspot wafer. Test indicated 2 small hotspots at 7:00. Cleaned the E-chuck with tooling and retested- OK. Performed IQC test. IQC test results were very close to results prior to shutdown. Checked EO-  OK SVG Dev: 54, SVG Dev2 55. System back up and ready to use.




More information about the asml-pcs mailing list